A 1.4 Times 10^-3 V/m electric field creates a 5.5
Answer #1

Drift velocity, v = I/ nA

V = (5.5 x 10^17)/(6 x 10^28 x pi x (0.75 x 10^-3)^2)

V = 5.19 x 10^-6 m/s

B)

Time, t = mv/eE

t = (9.11 x 10^-31 x 5.19 x 10^-6)/(1.6 x 10^-19 x 0.0014) = 2.11 x 10^-14 sec

Comment in case any doubt please rate my answer....